OpenStudy (anonymous):

A customer has an offer for 5 portrait sheets for 9.99 with additional portrait sheets priced at 19.99 each. If the customer desires to purchase 10 portrait sheets, how much would she save if she instead purchases a 10 portrait sheet package for $99.99? $19.99 Not Here $19.95 $9.99 $9.95

OpenStudy (anonymous):

not here right

OpenStudy (anonymous):

Can't understand your question

OpenStudy (anonymous):

1st option costs9.99 +5*19.99=9.99+99.95=119.94 or second option costs99.99 1st option minus second option is 119.94-99.99=19.95 So 19.95